OpenStudy (anonymous):

can you help me with my algebra homework please. 1. 5x-3y-15=0 2. x+2y+3=0 3.-3x+5y-15=0 4. y=6-6x

OpenStudy (anonymous):

Do you have to factor it?

OpenStudy (anonymous):

its about x and y intercepts

OpenStudy (anonymous):

What do the directions say?

OpenStudy (anonymous):

just find the x and y intercept for each answer..

OpenStudy (anonymous):

Hmmmm....

OpenStudy (anonymous):

could you hold on a couple minutes?

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

Ok so for the y intercept: you have 5x-3y-15=0 So you have to put that in y=mx+b form so subtract 5x from each side and add 15 to each side -5x+15=-3y divide each side by -3 5/3x-5=y

OpenStudy (anonymous):

In a y=mx+b equation the y intercebt is the b.... so in this case the y intercept is -5

OpenStudy (anonymous):

ok what about the x and could you just tell me the answers for all of them causse i know how to them..... but could u just please tell me the answers

OpenStudy (anonymous):

thanksssss :)

OpenStudy (anonymous):

yeah one second

OpenStudy (anonymous):

to find the x intercept you do the same thing but instead of equaling y it has to equal x

OpenStudy (anonymous):

so take the equation you have created: y=5/3x-5

OpenStudy (anonymous):

and make it equal x.... what you do is add five to each side y+5=5/3x now multiply each side by the reciprocal of the fraction (3/5) now you have: 3/5y+3=x

OpenStudy (anonymous):

the x intercept is 3

OpenStudy (anonymous):

And I can't do your homework for you, but if you'd like, tell me the answers first and after that I'll give you the correct answers.... that way you could see if you were right

OpenStudy (anonymous):

...and get some good practice